OpenStudy (anonymous):

The height of a side on a pyramid is 11 inches long, and its square base has sides that are 6 inches long. What is the surface area? 233 in.2 168 in.2 144 in.2 207 in.2

OpenStudy (anonymous):

total surface area = (area of square base) + (area of 4 triangles)

OpenStudy (anonymous):

Because you say the height of a side (instead of the height of the pyramid itself) then the answer is\[6 \times 6 (base) + 4 \times (11 \times 6 \div 2) = 168\]